OpenStudy (anonymous):

Find the values of x for which the series converges (write the answer in interval form) Sum of ((-6)^n)(x^n) starting from n=1 to infinity.

OpenStudy (dumbcow):

For a series of the form: \[\sum_{n=1}^{\infty} a^n\] it converges only if -1 < a < 1

OpenStudy (dumbcow):

\[-1 < -6x < 1\] \[-\frac{1}{6} <x < \frac{1}{6}\]

OpenStudy (dumbcow):

\[x \ne 0\]

OpenStudy (anonymous):

Thank you, this really makes sense! But how would I answer a question like: Find the sum of the series for those values of x. when this is an interval answer?